# Non-unital noise in a superconducting quantum computer as a computational resource for reservoir computing

### v2

@@ Results
By tuning the intensity of pure amplitude damping in simulations, the learning capabilities of the network are maximized for $\gamma\sim 0.02$. corresponding to which the echo state performs best for any task. This suggests the existence of an optimal critical regime for the network’s expressivity. The emergent properties of a quantum circuit depend on the rate of mid-circuit measurements performed on it [43]. This turns out to be the case also for the expressivity and the short-term memory capacity of the circuit acting as the reservoir. Indeed, when performing a complete measurement of the qubits in the register, the beneficial effect of amplitude damping appears stable under increasing noise intensity, at least for small values of $\lambda$ (approximately, $\gamma<0.1$). As reported in Fig. 2, this optimal value of $\gamma$ is followed by a plateau of suboptimal values, for which the performance remains acceptable. In the case of partial measurements of the qubit register, while the optimal noise regime persists for the same noise intensity with no deterioration in memory capacity or predictive accuracy, the plateau of suboptimal values is immediately disrupted by increasing the noise intensity. Moreover, the best performances of the network essentially do not depend on the number of measured qubit.
@@
We confirm our findings by experimenting on the IBM-Brisbane superconducting quantum computer, guided by the results of the noisy simulations. First, we employ the complete measurement of the register in order to enhance the stability under different noise intensities. As a further optimization, we introduce a tunable delay in executing rotations, deploying them as $RZ(\alpha)=RZ(\alpha/d)\circ\dots\circ RZ(\alpha/d)$ and the same for $RX$, where $\alpha$ denotes the angle of the desired rotation. Then, assuming that a delayed execution enlarges the amount of noise affecting the circuit evolution, we optimize the delay parameter $d$ by employing the emulator, before running the experiment on quantum hardware. However, there is a trade-off to consider. Although using multiple rotations instead of a single rotation on a quantum computer helps us achieve the desired increase in error, it also comes with a downside. Quantum computers have a limited allowance for the number of operations that can be submitted, which constrains our approach. We found an optimal compromise between performances and depth of the circuit for $d=4$. The experiment is conducted 5 times and the average results are shown with error bars of 1 standard deviation in Fig. 3. The experiment shows that the learning capabilities are considerably better when compared to the ideal execution without noise performed on the simulator. However, the NARMA-8 task clearly shows that when processing longer information windows is required, the unital component of the intrinsic noise in the quantum computer dominates, leading to a progressive deterioration in performance. This highlights the need to implement appropriate error mitigation techniques to eliminate the effects of unital noise or to adopt suitable strategies to amplify the impact of non-unital noise, which is undoubtedly recognized as essential in simulations.
@@ Discussion
We showed the beneficial effect of nonunital noise in quantum reservoir computing for time-dependent tasks. In particular, we identify the noise models characterized by non-unital channels as essential for guaranteeing a drastic improvement in the learning capabilities of a quantum echo state network that employs a register of superconducting qubits as the reservoir. Indeed, we observe in numerical simulations that amplitude damping ensures great accuracy in benchmark time-dependent tasks. The learning is confirmed by conducting the experiment on the superconducting computer. Nonunitality of the evolution of the quantum circuit has already been recognized as a key resource for reservoir computing [34]. Here we show that it can be naturally induced by employing intrinsic noise in the system. Moreover, the network’s response to noise qualitatively varies depending on whether complete or partial measurements are performed. In particular, in the case of complete measurements, the learning capabilities of the network remain effective over a sufficiently broad range of noise intensity values, confidently encompassing the amount of noise present in current quantum hardware. These facts pave the way for use-case results in real-world applications. Moreover, from the perspective of using quantum hardware, our results offer valuable insights that can inform best practices for developing and applying suitable error mitigation techniques. On the one hand, our observations suggest prioritizing the mitigation of dephasing and readout errors for the application of quantum reservoir computing protocols while ignoring the correction of dissipation effects. Moreover, partially effective error mitigation techniques may be employed to reduce the number of measurements since the optimal performances of the network persist in the case of partial measurement of the circuit. As a consequence, employing error control algorithms, that allow positioning within this regime, can lead to significant savings in memory consumption and execution time on quantum hardware by reducing the total number of mid-circuit measurements. Our findings are supported by the universality approximation property as proved in the Supplementary Material, which is fulfilled by our gate-based echo state network under the effect of a general set of noise quantum channels, including the amplitude damping channel.
@@ Methods
Our quantum echo state network exploits a $N$-qubits quantum register to encode and process time-dependent information and reproduce an input-output nonlinear mapping $S(u)=\hat{y}$ between two time-dependent sequences $u=\left\{u_{t}\right\}_{t=0,\dots,L}$ and $\hat{y}=\left\{\hat{y}_{t}\right\}_{t=0,\dots,L}$. We refer to $u$ and $\hat{y}$ as the input and the target sequence, respectively.
Denoting with $\rho_{t}$ the density operator describing the $N$-qubits system at time $t$, its temporal evolution $\rho_{t+1}=\mathcal{T}_{u_{t+1}}(\rho_{t})$ is the composition of a noisy quantum channel, including the unitary gates that encode the input signal $u$, and a mid-circuit measurement operation. Precisely, at each time step, the system first evolves with the quantum evolution
@@
In Equation 5, $U(u_{t+1})$ represents some unitary gates implemented by the quantum computer. Our model includes CX operators and rotations, whose angles are dependent on the incoming input value. Precisely, we encode the input value in the angle of unitary rotations as $RX(u_{t})$ and $RZ(u_{t})$. The architecture adopted in our work is described in Fig 1. $\mathcal{E}$ is a quantum channel that models the noise in the quantum register. Then, a mid-circuit measurement of the register is performed and the system is let evolve in the collapsed state. Formally, we have
@@
with, denoting with $m_{s,t+1}\in\left\{0,1\right\}$ the outcome of the measure of the qubits $s$ at time $t+1$,
@@
Here, $M$ is the number of measured qubits in the register at each time step. We remark that both the case of a complete measurement, $M=N$, and the case of a partial measurement, $M<N$, are treated in the Results. The measurement results are taken as the reservoir signal passes throughout the processing. Explicitly, the reservoir signal at time $t$ is the expectation value of the $\sigma_{Z(i)}$ Pauli operator of each qubit, taken separately,
@@
In our numerical experiments, the expectation value of each $\sigma_{Z(i)}$ is obtained by repeating the experiment and computing $\langle\sigma_{Z(i)}\rangle_{\rho_{t}}=\frac{1}{S}\left(\sum 1_{m_{i,t}=0}-\sum 1_{m_{i,t}=1}\right),$ where the sum is taken over all shots.
@@
Reservoir computing aims to reproduce best a given nonlinear map $S(u)=\hat{y}$ by training the linear readout $h$. We will give examples of such learning tasks $S(u)$ later in the work. After an initial washout period $T_{\text{wo}}$, required to erase the dependence on the initial state of the network, the reservoir signal is extracted until the end of the training interval $T_{\text{tr}}$ and it is stored in the matrix $H=\left\{z_{t}\right\}_{t=T_{\text{wo}}+1,\dots,T_{\text{tr}}}$. After the quantum evolution, the optimal readout weights are computed. Precisely, the weights of the linear readout $y=Hw$ are chosen as the set of parameters $w$ that minimizes the distance $\norm{\hat{y}-y}$. It can be easily computed by linear regression or exploiting the Moore-Penrose pseudoinverse matrix, namely $w=(H^{T}H)^{-1}H^{T}\hat{y}_{\text{tr}}$, where $\hat{y}_{\text{tr}}=\left\{\hat{y}_{t}\right\}_{t=T_{\text{wo}}+1,\dots,T_{\text{tr}}}$ is the portion of target sequence used for the training. We refer to Ref. [30] for further details on the training algorithm. After the training, the predicted values of $y$ are given by
@@
It is worthwhile to notice that, thanks to the intrinsic multitasking nature of reservoir computing, the same reservoir signal can be exploited to reproduce different maps, only repeating the training to compute the optimal weights. This fact allows for the reuse of information stored in the reservoir for different tasks that share the same input, notably reducing execution time.
@@
Our experiments aim to demonstrate the beneficial effect of intrinsic noise in a superconducting quantum computer for reservoir computing. First, to verify whether quantum noise improves the fading memory, we study its effect on the correlation among different measurement outcomes of the same qubit over time. Then, the performances of our computational architecture under the effect of quantum noise are analyzed by exploiting standard benchmarks for recurrent neural networks. We measure the short-term memory capacity of our network using the memory capacity metric, which quantifies the amount of variance in the delayed input that can be recovered from the trained output [9]. To compute it, the system is required to replicate an input sequence $\left\{u_{t}\right\}_{t=0}^{T}$ delayed by a time interval $d$. Namely, the system is required to approximate the functional $\hat{y}_{t}=u_{t-d}$. Then, the $d$-memory capacity is defined as
$\mathrm{MC}_{d}=\frac{\mathrm{Cov}\left(y,\hat{y}\right)}{\mathrm{Var}(y)\mathrm{Var}(\hat{y})}\,,$ (10)
@@
normalized so that $\mathrm{MC}\in[0,1]$. We remark that $\mathrm{MC}$ values closer to 1 indicate a higher memory capacity of the system. We fix $d_{\text{max}}=10$ in the numerical experiments.
The predictive capabilities of our echo state network are tested with the so-called NARMA task. It is a nonlinear filter with past dependence, commonly used as a benchmark for the computational power of time-dependent learning systems [44]. Precisely, the NARMA-$p$ task is formally defined as
$\hat{y}_{t}=0.4y_{t}+0.1\left(\sum_{l=0}^{p-1}y_{t-l}\right)+u_{t-p+1}+0.1\,.$ (12)
We pick $p=2,5\text{ and }8$ in the paper. The accuracy of the prediction is quantified with the normalized mean-square error, expressed as
@@
The effect of noise in a quantum computer can be simulated by repeatedly applying a quantum channel to the density operator that describes the state of the qubit register. Precisely, in numerical simulations, a quantum channel $\mathcal{E}$, representing a certain noise model, is applied to each qubit after the action of each unitary gate $U$, as shown in Equation  5. First, we investigate the effect of three typical ideal models of quantum noise. Namely, we test the effect of amplitude damping, phase damping, and depolarizing noises. We recall that a quantum channel can always be expressed in terms of its Kraus decomposition,
@@
We refer to Table 1 in the Supplementary Material for a detailed description of the Kraus operators of the three noise models considered in this work. The noise intensity is tuned by a parameter appearing in the Kraus decomposition, which we indicate with $\gamma$ throughout the paper. Smaller $\gamma$ values correspond to a final quantum state less affected by noise. After the preliminary analysis of the effect of an ideal pure noise, we test our architecture with a more realistic model of quantum noise, which faithfully emulates noise in a superconducting device. To do this, we apply the quantum channels that model the noise of IBMQ backends freely available to the user. The simulation of a backend noise model implemented by Qiskit [45] exploits a combination of thermal relaxation and depolarizing channels.
@@
In this work, we fix $N=7$ as the number of qubits. In numerical simulations, we consider a total timespan of $L=200$, with a washout period of 10 time steps ($T_{\text{wo}}=10$), while the test interval is 60 time steps long. Each circuit is run for $S=10^{5}$ shots for a precise reconstruction of the average value of observables. All the simulations with an ideal noise model are run over 100 different values of $\gamma$. Preliminary experiments are conducted using the Qiskit simulator [45]. In experiments involving the quantum hardware, we significantly reduce the depth of the circuit due to technical limitations of the cloud service employing the IBM quantum service and cost management. In particular, we fix $L=80$, and for each experiment, we submit 90 jobs in batch, with 5 circuits each and 30 shots per shot, for a total of 13500 shots. We used the IBM_BRISBANE quantum processing unit, which mounts an Eagle processor, equipped with 127 qubits. For the simulations, we used a 36-core Intel Xeon processor (4.3 GHz) with 128 GB of RAM.

### v3

@@ Results
By tuning the intensity of pure amplitude damping in simulations, the learning capabilities of the network are maximized for $\gamma\sim 0.02$. corresponding to which the echo state performs best for any task. This suggests the existence of an optimal critical regime for the network’s expressivity. The emergent properties of a quantum circuit depend on the rate of mid-circuit measurements performed on it [50]. This turns out to be the case also for the expressivity and the short-term memory capacity of the circuit acting as the reservoir. Indeed, when performing a complete measurement of the qubits in the register, the effect of amplitude damping appears stable under increasing noise intensity, at least for small values of $\lambda$ (approximately, $\gamma<0.1$). As reported in Fig. 3, this optimal value of $\gamma$ is followed by a plateau of suboptimal values, for which the performance remains acceptable. In the case of partial measurements of the qubit register, while the optimal noise regime persists for the same noise intensity with no deterioration in memory capacity or predictive accuracy, the plateau of suboptimal values is immediately disrupted by increasing the noise intensity. Moreover, the best performances of the network essentially do not depend on the number of measured qubits.
@@
We confirm our findings by experimenting on the IBM-Brisbane superconducting quantum computer, guided by the results of the noisy simulations. First, we employ the complete measurement of the register in order to improve stability under various noise intensities. As a further optimization, we introduce a tunable delay in executing rotations, deploying them as $RZ(\alpha)=RZ(\alpha/d)\circ\dots\circ RZ(\alpha/d)$ and the same for $RX$, where $\alpha$ denotes the angle of the desired rotation. Then, assuming that a delayed execution enlarges the amount of noise affecting the circuit evolution, we optimize the delay parameter $d$ by employing the emulator, before running the experiment on quantum hardware. However, there is a trade-off to consider. Although using multiple rotations instead of a single rotation on a quantum computer helps us achieve the desired increase in error, it also comes with a downside. Quantum computers have a limited allowance for the number of operations that can be submitted, which constrains our approach. We found an optimal compromise between performance and depth of the circuit for $d=4$. The experiment is conducted 5 times and the average results are shown with error bars of 1 standard deviation in Fig. 4. The experiment shows that the learning capabilities are considerably better when compared to the ideal execution without noise performed on the simulator. However, the NARMA-8 task clearly shows that when processing longer information windows is required, the unital component of the intrinsic noise in the quantum computer dominates, leading to a progressive deterioration in performance. This highlights the need to implement appropriate error mitigation techniques to eliminate the effects of unital noise or to adopt suitable strategies to amplify the impact of non-unital noise, which is undoubtedly recognized as essential in simulations.
@@ Discussion
We demonstrate the beneficial effect of damping in gate-based quantum reservoir computing for time-dependent tasks. Firstly, we showed that dissipation is essential for the functioning of a gate-based echo state network. Indeed, the repeated measurements needed for information decoding naturally steer the internal state of the circuit towards the maximally mixed state, making the input encoding ineffective after a short time window. In this work, we show that non-unital noise, as amplitude damping, prevents this phenomenon, leading to a long-term effective reservoir computation. Moreover, as a further theoretical support to empirical findings, we identify the properties of the quantum dynamics of the reservoir that ensure the universal approximation property of the reservoir computer. Indeed, they turn out to be the same properties ensuring the functioning of the reservoir computer. Based on our theoretical finding, we confirm early empirical results [31, 39] by observing through numerical simulations that amplitude damping ensures great accuracy in benchmark time-dependent tasks. The learning is confirmed by conducting the experiment on the superconducting computer. Non-unitality of the evolution of the quantum circuit has already been recognized as a key resource for reservoir computing [37]. Here we show that it can be naturally induced by employing intrinsic noise in the system. Moreover, the network’s response to noise qualitatively varies depending on whether complete or partial measurements are performed. In particular, in the case of complete measurements, the learning capabilities of the network remain effective over a sufficiently broad range of noise intensity values, confidently encompassing the amount of noise present in current quantum hardware. These facts pave the way for use-case results in real-world applications. Moreover, from the perspective of using quantum hardware, our results offer valuable insights that can inform best practices for developing and applying suitable error mitigation techniques. On the one hand, our observations suggest prioritizing the mitigation of dephasing and readout errors for the application of quantum reservoir computing protocols while ignoring the correction of dissipation effects. Partially effective error mitigation techniques may be employed to reduce the number of measurements since the optimal performances of the network persist in the case of partial measurement of the circuit. As a consequence, employing error control algorithms that allow positioning within this regime can lead to significant savings in memory consumption and execution time on quantum hardware by reducing the total number of mid-circuit measurements. Nevertheless, the relatively poor experimental results are due to the fact that current quantum computers already have error rates below the noise regime that ensures the functioning of an echo state network gate-based with mid-circuit measurements. For this reason, new algorithms that increase qubit damping are needed for effective implementation.
@@ Methods
Our gate-based echo state network employs an N-qubit circuit to encode and process time-dependent information. Then, a linear readout is trained to replicate a mapping $S(u)=\hat{y}$ between an input sequence $u=\left\{u_{t}\right\}_{t=0,\dots,L}$ and a target sequence $\hat{y}=\left\{\hat{y}_{t}\right\}_{t=0,\dots,L}$. In each execution, the time series is entirely processed by executing a single circuit that consists of unitary evolution, which is used for encoding the input value, and mid-circuit measurements of the qubits, whose outcomes are used for computing the expected value of the observables employed as readout. In particular, after the encoding of an input value, the mid-circuit measurement of the register is performed on the computational basis and, as opposed to other approaches [37, 51], the system is let evolve in the collapsed state after the measurements. The entire circuit, which encodes progressively the input time series, is executed for $S$ shots for computing the deterministic expected value of the observables used as readout. We describe in the following the one-step evolution of the systems, which is replicated for each value of the input time series. We denote with $\rho_{t}$ the density that describes the state of the N-qubit register at time-step $t$. We remark that the density matrix $\rho_{t}$ depends formally on any value of the input $u_{t^{\prime}}$ with $t^{\prime}\leq t$, when considered as a statistical mixture of the $S$ repetitions of the circuit, that are conditioned on the outcome of the mid-circuit measurements. Indeed, the outcome probabilities directly depend on the value of the input. Firstly, $\rho_{t}$ is subject to a unitary evolution affected by some noise, modelled as a quantum channel applied after the unitary evolution, namely
@@
Here, $\mathcal{E}$ denotes the quantum channel that models the quantum noise, described later in the paper. The input $u_{t+1}$ is encoded in the angles of some rotations which constitute the unitary gate $U$. The circuit, described in detail in Fig 1, is composed of rotations RX, RY, and CNOT gates, which entangle pairs of qubits at a time. After the input injection, a mid-circuit measurement of the qubits is performed, and the system is let evolve in the collapsed state. Thus, for each execution of the circuit, the evolution is conditioned by the measurement outcomes. This mechanism is key to integrating past information into the system, enabling the processing of time-dependent data without requiring hidden qubits for memory retention, in contrast to the approach used in Ref. [37]. In particular, we assume that each qubit is measured separately and simultaneously, so that each measurement does not affect the outcome of the others. Formally, we write the effect of the measurement of $M$ out of $N$ qubits as
@@
where
@@
denoting $m_{s,t+1}\in\left\{0,1\right\}$ the outcome of the measurement of the qubits $s$ at time $t+1$. In particular, we measure either all the qubits employed in the circuits ($M=N$) or a smaller subset ($M<N$). In both cases, the reservoir signal at time-step $t$ is constructed online as the expectation value of the $Z_{i}$ Pauli operators, corresponding to the measured qubits, namely
@@
Explitcly, we construct the expectation value of the observable $\sigma_{Z(i)}$ by collecting the $S$ outcomes of each measurement of the $i^{th}$ qubit, and then computing
$\langle\sigma_{Z(i)}\rangle_{\rho_{t}}=\frac{1}{S}\left(\sum 1_{m_{i,t}=0}-\sum 1_{m_{i,t}=1}\right),$ (12)
where $1_{m_{i,t}=k}$ assumes value $1$ when $m_{i,t}=k$ and 00 otherwise, with $k=0,1$, since indeed the eigenvalues of the observable $\sigma_{Z}$ are $1$ and $-1$, corresponding to the eigenstate |$\ket{0}$ and |$\ket{1}$, respectively. This procedure amounts to computing the expected value of the global observables $\mathbb{I}^{\otimes i-1}\otimes\sigma_{Z(i)}\otimes\mathbb{I}^{\otimes N-i}$. This is consistent with the assumption about the effect of local measurements since these observables commute.
@@
Reservoir computing aims to best reproduce a given nonlinear map $S(u)=\hat{y}$ by training the linear readout $h$. We will give examples of such learning tasks $S(u)$ later in the work. After an initial washout period $T_{\text{wo}}$, required to erase the dependence on the initial state of the network, the reservoir signal is extracted until the end of the training interval $T_{\text{tr}}$ and it is stored in the matrix $H=\left\{z_{t}\right\}_{t=T_{\text{wo}}+1,\dots,T_{\text{tr}}}$. After the quantum evolution, the optimal readout weights are computed. Precisely, the weights of the linear readout $y=Hw$ are chosen as the set of parameters $w$ that minimizes the distance $\norm{\hat{y}-y}$. It can be easily computed by linear regression or exploiting the Moore-Penrose pseudoinverse matrix, namely $w=(H^{T}H)^{-1}H^{T}\hat{y}_{\text{tr}}$, where $\hat{y}_{\text{tr}}=\left\{\hat{y}_{t}\right\}_{t=T_{\text{wo}}+1,\dots,T_{\text{tr}}}$ is the portion of target sequence used for the training. We refer to Ref. [30] for further details on the training algorithm. After the training, the predicted values of $y$ are given by
@@
It is worth mentioning that, thanks to the intrinsic multitasking nature of reservoir computing, the same reservoir signal can be exploited to reproduce different maps, only repeating the training to compute the optimal weights. This fact allows for the reuse of information stored in the reservoir for tasks that share the same input, notably reducing execution time.
@@
Noise model Description Kraus decomposition
Our experiments aim to demonstrate the beneficial effect of intrinsic noise in a superconducting quantum computer for reservoir computing. First, to verify whether quantum noise improves the fading memory, we study its effect on the correlation among several measurement outcomes of the same qubit over time. Then, the performances of our computational architecture under the effect of quantum noise are analyzed by exploiting standard benchmarks for recurrent neural networks. We measure the short-term memory capacity of our network using the memory capacity metric, which quantifies the amount of variance in the delayed input that can be recovered from the trained output [9]. To compute it, the system is required to replicate an input sequence $\left\{u_{t}\right\}_{t=0}^{T}$ delayed by a time interval $d$. Namely, the system is required to approximate the functional $\hat{y}_{t}=u_{t-d}$. Then, the $d$-memory capacity is defined as
$\mathrm{MC}_{d}=\frac{\mathrm{Cov}^{2}\left(y,\hat{y}\right)}{\mathrm{Var}(y)\mathrm{Var}(\hat{y})}\,,$ (14)
@@
normalized so that $\mathrm{MC}\in[0,1]$. We remark that $\mathrm{MC}$ values closer to 1 indicate a higher memory capacity of the system. We fix $d_{\text{max}}=10$ in the numerical experiments. This definition of memory capacity should not be confused with the one introduced in [52]. The latter also considers the capacity of non-linear memory, whereas we consider only linear short-term memory. Our choice is driven by a more practical and task-specific score for benchmarking. In addition, the total capacity of [52] is upper-bounded by the number of internal states, whereas our memory capacity depends on $d_{\text{max}}$.
The reconstruction capabilities of our echo state network are tested with the so-called NARMA task. It is a nonlinear filter with past dependence, commonly used as a benchmark for the computational power of time-dependent learning systems [53]. Precisely, the NARMA-$p$ task is formally defined as
$\hat{y}_{t}=0.4\,y_{t-1}+0.1\,\left(\sum_{l=0}^{p-1}y_{t-l}\right)+\,u_{t-p+1}u_{t}+0.1\,.$ (16)
We pick $p=2,5\text{ and }8$ in the paper. The accuracy of the reconstruction is quantified with the normalized mean-square error, expressed as
@@
The effect of noise in a quantum computer can be simulated by repeatedly applying a quantum channel to the density operator that describes the state of the qubit register. Precisely, in numerical simulations, a quantum channel $\mathcal{E}$, representing a certain noise model, is applied to each qubit after the action of each unitary gate $U$, as shown in Equation  8. First, we investigate the effect of three typical ideal models of quantum noise. Namely, we test the effect of amplitude damping, phase damping, and depolarizing noises. We recall that a quantum channel can always be expressed in terms of its Kraus decomposition,
@@
We refer to Table 2 for a detailed description of the Kraus operators of the three noise models considered in this work. The noise intensity is tuned by a parameter appearing in the Kraus decomposition, which we indicate with $\gamma$ throughout the paper. Smaller $\gamma$ values correspond to a final quantum state less affected by noise. After the preliminary analysis of the effect of an ideal pure noise, we test our architecture with a more realistic model of quantum noise, which faithfully emulates noise in a superconducting device. To do this, we apply the quantum channels that model the noise of IBMQ backends freely available to the user. The simulation of a backend noise model implemented by Qiskit [54] exploits a combination of thermal relaxation and depolarizing channels.
@@
In this work, we fix $N=7$ as the number of qubits. In numerical simulations, we consider a total timespan of $L=200$, with a washout period of 10 time steps ($T_{\text{wo}}=10$), while the test interval is 60 time steps long. Each circuit is run for $S=10^{5}$ shots for a precise reconstruction of the average value of observables. All simulations with an ideal noise model are run over 100 values of $\gamma$. Preliminary experiments are conducted using the Qiskit simulator [54]. In experiments involving quantum hardware, we significantly reduce the circuit depth according to the limitations of the cloud service employing the IBM quantum service and cost management. In particular, we fix $L=80$, and for each experiment, we submit 90 jobs in batch, with 5 circuits each and 30 shots per circuit, for a total of 13500 shots. We used the IBM_BRISBANE quantum processing unit, which mounts an Eagle processor, equipped with 127 qubits. For the simulations, we used a 36-core Intel Xeon processor (4.3 GHz) with 128 GB of RAM.